you arrive at another model rocket competition and you have remembered to bring your small skid fuel engines. your rocket is launched form the rest and the solid fuel engine delivers a constant acceleration of 8.2 m/s^2 for 5 seconds, after which the fuel is used up. Using homer's formula, S=1/2at^2, find the altitude of the rocket when the engine shuts down.

Answer by Alan3354(69443)

8.2 m/s^2 for 5 seconds,
--------
s = at^2/2 = 8.2*25/2
= 102.5 m